# Carrer Pic de Peguera 13, 17003 Girona, Spain

### v1

@@ 5.2 Data Encoding and Augmentation
Traditionally, CNN-based approaches (Zhu et al., 2018b, 2019; Wan et al., 2022) require the EBC output to be converted from raw streams of events data to 2-channel event images, where one channel represents positive polarities and the other the negative polarities. To support this, eWiz provides different encoders. Currently, two encoding schemes are implemented: Gaussian (Ding et al., 2022) and event counting Nguyen et al. (2019). The former computes pixel values by summing each event’s polarity weighted by a normalized Gaussian kernel based on the event timestamp and a scaling factor $\lambda$ for each bin. Whereas, the latter simply groups events into fixed time intervals, with pixel values computed by summing the polarities of all events within each bin.
@@ 5.4 Algorithms
eWiz not only supports data processing for NN-based approaches but also algorithms, such as contrast maximization (Shiba et al., 2022), that directly operate on raw event streams. Contrast maximization involves warping the event stream based on a proposed optical flow, iteratively optimizing the optical flow field to maximize the contrast computed between the warped and the original event streams. Higher contrast, indicated by well-defined edges, signals improved alignment of events. As such, eWiz also provides warping tools that allow the implementation of various motion models based on the camera movement type. In the long run, eWiz will support not only optimization-based algorithms but also deep learning-based and SNN-based alternatives. We plan to add more functionalities, such as ground truth generation via optimization and SNN-based training algorithms.
@@ 5.5 Loss Functions and Evaluation Metrics
eWiz supports multiple loss functions for training CNNs and SNNs, currently focused on optical flow prediction. For instance, it supports the photometric loss (Zhu et al., 2018b), as a self-supervised approach to minimize pixel-wise intensity differences between two grayscale images encompassing an event stream. The predicted flow field is used to warp the grayscale images and optimize alignment. This loss can be combined with Charbonier loss (Sun et al., 2014), a smooth approximation of the L1 loss that is more sensitive to large outliers, as well as a smoothness loss serving as a regularizer to address the aperture problem and ensure continuous flow across sparse event-based data. In addition, the library includes losses specifically designed for the contrast maximization algorithm (Gallego et al., 2018, 2019) to evaluate warped event streams. These include image variance and gradient magnitude loss functions, along with their improved variants presented by Shiba et al. (2022), where the losses are normalized and applied in a multi-focal manner. To further enhance performance, eWiz also provides a regularizer that smooths the predicted flow across the image, helping address the sparsity of event-based data and ensuring continuous flow, ultimately improving convergence during training.
Moreover, eWiz can be used not only to evaluate the performance of trained models but also to track their performance during training. Currently, the library supports performance metrics for optical flow prediction tasks, such as the Average Endpoint Error (AEE) (Zhu et al., 2018b), which is defined as the distance between the endpoints of the predicted and the ground truth flow vectors, and the Average Angular Error (AAE) (Zhu et al., 2018b), defined as the angular difference between the prediction and ground truth flows, both averaged over the number of pixels. Due to the sparsity of event-based data, these metrics are only computed for pixels where events have occurred. Another implemented metric is the outlier percentage, which refers to the percentage of pixels in the image with an endpoint error higher than a certain value. We define XPE the overall outlier percentage of pixels with endpoint errors greater than X.
@@ 6 Dataset Structure
The eStonefish-scenes dataset follows the eWiz format, which stores data in HDF5 files (The HDF Group, 2006). Specifically, events.hdf5 holds the event streams, gray.hdf5 contains grayscale image sequences associated with each event stream, and flow.hdf5 stores the optical flow information. Metadata and configuration details, including sensor specifications and recording parameters, are saved in props.json. This structure facilitates straightforward access to each data type, improving data handling and compatibility with various processing pipelines. The directory structure is summarized below:
[…]
Figure 6: Structure of the data repository.
The official documentation provides detailed information on the contents of each HDF5 file. However, users can easily access or write data through the readers and writers included in eWiz without needing to handle these details directly.
@@ 7 Conclusion
This work addresses the challenge of developing reliable optical flow prediction networks for underwater environments, where collecting large-scale, annotated, event-based optical flow datasets is exceptionally difficult due to the complexities of underwater imaging. To overcome this, this work introduces the eStonefish-scenes dataset, along with a user-friendly pipeline to extend and adapt the dataset. To the best of our knowledge, no prior event-based datasets exist for underwater applications. The dataset incorporates scenes capturing event-based data from an ROV navigating a dynamic, coral-rich environment filled with diverse marine flora and fauna. We also introduce Stonefish-scenegen, an automated scene generator that eliminates the need for manually creating time-consuming reef environments, and Stonefish-boids, a package that simulates realistic schooling behaviour of fish. The dataset and supporting packages were created with the goal of accelerating research on EBCs for underwater applications, especially for training CNNs and SNNs. We also present eWiz, a comprehensive event-based data manipulation library that includes an optimized event data format. The library features data converters, writers, readers, and processing modules focused on data management, compression, and optimization. Additionally, it offers various loss functions, performance metrics, and visualizers to facilitate work with these datasets, along with an implementation of the contrast maximization algorithm. Both eWiz and eStonefish-scenes are designed to streamline research on EBCs and motion estimation for autonomous underwater vehicles by providing easy-to-use libraries for generating, processing, and evaluating event-based datasets.
@@
Jad Mansour was supported by the Joan Oró Grant no. 2024 FI-2 00762. The study was also supported in part by the Spanish government through the SIREC project no. PID2020-116736RB-IOO.
@@
Jad Mansour: Conceptualization, Methodology, Software, Data Curation, Writing - original draft; Sebastian Realpe: Software, Writing - original draft; Hayat Rajani: Conceptualization, Methodology, Supervision, Validation, Writing - original draft; Michele Grimaldi: Software, Writing - review & editing; Rafael Garcia: Project Administration, Supervision, Writing - review & editing; Nuno Gracias: Project Administration, Supervision, Writing - review & editing.

### v2

@@ 5.2 Data Encoding and Augmentation
Traditionally, CNN-based approaches (evflownetZhu18; egoflownetZihao19; biflowWan22) require the EBC output to be converted from raw streams of event data to two-channel event images, where one channel represents positive polarities and the other represents negative polarities. To support this, eWiz provides different encoders. Currently, two encoding schemes are implemented: Gaussian (gaussianDing22) and event counting (countNguyen19). The former computes pixel values by summing each event’s polarity weighted by a normalized Gaussian kernel based on the event timestamp and a scaling factor $\lambda$ for each bin. The latter simply groups events into fixed time intervals, with pixel values computed by summing the polarities of all events within each bin.
@@ 5.4 Algorithms
eWiz not only supports data processing for NN-based approaches but also supports algorithms, such as contrast maximization (mcShiba2022), that operate directly on raw event streams. Contrast maximization involves warping the event stream based on a proposed optical flow, iteratively optimizing the optical flow field to maximize the contrast computed between the warped and original event streams. Higher contrast, indicated by well-defined edges, signals improved alignment of events. As such, eWiz also provides warping tools that enable the implementation of various motion models based on the camera movement type.
@@ 5.5 Loss Functions and Evaluation Metrics
eWiz supports multiple loss functions for training CNNs and SNNs. For instance, it supports the photometric loss (evflownetZhu18) as a self-supervised approach to minimize pixel-wise intensity differences between two grayscale images captured immediately before and after the input events stream. The predicted flow field is used to warp the grayscale images and optimize alignment. This loss can be combined with Charbonier loss (charbonnierSun14), a smooth approximation of the L1 loss that is less sensitive to outliers, as well as a smoothness loss serving as a regularizer to address the aperture problem and ensure continuous flow across sparse event-based data. In addition, the library includes losses specifically designed for the contrast maximization algorithm (mcGallego2018; mcGallego2019) to evaluate warped event streams. These include image variance and gradient magnitude loss functions, along with the improved variants presented by mcShiba2022, where the losses are normalized and applied in a multi-focal manner. To further enhance performance, eWiz also provides a regularizer that smooths the predicted flow across the image, helping address the sparsity of event-based data and ensuring continuous flow, ultimately improving convergence during training.
Moreover, eWiz can be used not only to evaluate the performance of trained models but also to track their performance during training. Currently, the library supports performance metrics for optical flow prediction tasks, such as the Average Endpoint Error (AEE) (evflownetZhu18). AEE is defined as the distance between the endpoints of the predicted and ground-truth flow vectors, while the AAE measures the angular difference between the predicted and ground-truth flows. Both metrics are averaged over the number of pixels. Due to the sparsity of event-based data, these metrics are only computed for pixels where events have occurred. Another implemented metric is the outlier percentage, which refers to the percentage of pixels in the image with an endpoint error exceeding a specified threshold.
@@ 6 Dataset Structure
The eStonefish-Scenes dataset follows the eWiz format, which stores data in HDF5 files (hdf5Format). Specifically, events.hdf5 holds the event streams, gray.hdf5 contains grayscale image sequences associated with each event stream, and flow.hdf5 stores the optical flow information. Metadata and configuration details, including sensor specifications and recording parameters, are saved in props.json. This structure facilitates straightforward access to each data type, improving data handling and compatibility with various processing pipelines. The directory structure is summarized below:
dataset_root     events.hdf5     gray.hdf5     flow.hdf5     props.json
Figure 6: Directory structure of the eStonefish-Scenes data repository.
7 Ground Truth Acquisition
A crucial aspect of validating a synthetic dataset is evaluating the sim-to-real transferability (surveyPitkevich24). This is done in Section 8, using a real-world dataset collected under controlled conditions in an indoor testing pool. The dataset is described in this section.
7.1 Experimental Setup
Data was collected using a BlueROV2 equipped with a downward-facing DAVIS346 event camera. The ROV was deployed in the indoor testing pool at the Center for Research in Underwater Robotics and Vision (CIRS) in Girona. The ROV was manually controlled to perform six-degree-of-freedom motions, including translations along the forward–backward, lateral, and vertical axes, as well as yaw rotations.
The DAVIS346 sensor recorded both asynchronous event streams and grayscale images at a resolution of 346×260 pixels with a $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ field of view. Eight sequences were recorded, each lasting 20 seconds. The pool was illuminated with natural ambient light. A high-resolution reef-like poster covering the entire pool floor was used to provide underwater textures, with the camera positioned at an average distance of 2.5 meters from the poster surface.
Figure 7: CIRS testing pool with floor poster used for ground-truth acquisition.
7.2 Ground Truth Flow Computation
Following data acquisition, we estimated ground-truth optical flow using homography-based registration of the grayscale frames to the poster in the pool.
7.2.1 Camera Calibration
The DAVIS346 calibration was performed using a checkerboard pattern, in front of which the camera was moved at different orientations. The grayscale images were then undistorted using the resulting intrinsics. This correction mitigated distortion-induced geometric errors during image-to-poster registration. The resulting flow fields were then reprojected into the undistorted domain to maintain consistency with the recorded event-based data.
7.2.2 Flow Computation
For each undistorted grayscale frame, SIFT features were extracted and matched to a fixed reference image of the poster. Using matched keypoints $\{(p_{i}^{\text{frame}},p_{i}^{\text{poster}})\}$, we estimated a homography $H_{\text{image}\rightarrow\text{poster}}$ via RANSAC:
$p_{i}^{\text{poster}}\sim{}^{P}H_{i}\cdot p_{i}^{\text{image}}$ (1)
For two consecutive frames $A$ and $B$, the relative homography was computed as:
$^{B}H_{A}={}^{P}H_{B}^{-1}\cdot{}^{P}H_{A}$ (2)
To generate dense optical flow, we defined a grid of pixel coordinates $\{x_{i}\}$ in frame $A$, and warped them using the relative homography:
$\hat{x}_{i}=\frac{{}^{B}H_{A}\cdot[x_{i},y_{i},1]^{T}}{z_{i}},\quad f_{i}=\hat{x}_{i}-x_{i}$ (3)
This yields a dense flow field $f\in\mathbb{R}^{H\times W\times 2}$ between each pair of grayscale frames.
Figure 8: ROV inside the testing pool during data acquisition.
7.2.3 Outlier Rejection
During homography estimation, outlier rejection was performed using RANSAC. The inlier threshold was empirically set to 3 pixels based on the reprojection error between matched keypoints. For each frame, the proportion of inliers was computed as:
$\eta=\frac{N_{\text{inliers}}}{N_{\text{total}}}\times 100$ (4)
where $N_{\text{inliers}}$ denotes the number of keypoints satisfying the reprojection constraint:
$^{P}H_{i}p^{\text{image}}_{i}-p^{\text{poster}}_{i}||<3\text{px}$ (5)
The distribution of $\eta$ was analysed across all frames in each sequence to assess registration robustness. Histograms of these values indicated that, for the majority of frames, the outlier percentage $(100-\eta)$ remained below 30%, confirming the stability of the chosen threshold across the dataset.
7.3 Uncertainty Estimation
To estimate the uncertainty in the homography-based optical flow, we used a Monte Carlo method based on keypoint reprojection errors. For each frame, SIFT keypoints from the reference poster image were matched to the current frame, and a homography $H_{\text{poster}\rightarrow\text{image}}$ was estimated using RANSAC. The reprojection error was computed by projecting each poster keypoint into the image frame and comparing it with its corresponding detected keypoint:
$\epsilon_{i}=\left\|H_{\text{poster}\rightarrow\text{image}}\cdot p_{i}^{\text{poster}}-p_{i}^{\text{image}}\right\|$ (6)
The reprojection uncertainty for each frame was computed by aggregating all point-wise reprojection errors in both $x$ and $y$ directions into a single vector. Let the reprojection error for keypoint $i$ be $\boldsymbol{\epsilon}_{i}=(\epsilon_{x,i},\epsilon_{y,i})$. The overall noise level was estimated as the standard deviation of all concatenated components:
$\sigma=\sqrt{\frac{1}{2N-1}\sum_{i=1}^{N}\left[(\epsilon_{x,i}-\bar{\epsilon}_{x})^{2}+(\epsilon_{y,i}-\bar{\epsilon}_{y})^{2}\right]}$ (7)
where $N$ is the number of matched keypoints. The resulting $\sigma$ values were then averaged across all frames to obtain a representative uncertainty estimate for the dataset. This value was used as the noise scale in a Monte Carlo simulation.
For each of $N_{s}$ Monte Carlo iterations, we:
  1. 1.
    Added Gaussian noise to the poster keypoints:
    $\tilde{p}_{i}=p_{i}^{\text{poster}}+\mathcal{N}(0,\sigma^{2})$ (8)
  2. 2.
    Re-estimated the homography $\tilde{H}_{\text{poster}\rightarrow\text{image}}$ from noisy keypoints.
  3. 3.
    Recomputed the dense optical flow $f_{i}^{(n)}$ from the resulting relative homography.
Each pair of noisy homographies, $\tilde{H}_{a}$ and $\tilde{H}_{b}$, was used to warp a uniform grid of image coordinates and compute the corresponding optical flow sample $\mathbf{f}^{(n)}\in\mathbb{R}^{H\times W\times 2}$. Let $\mathbf{d}^{(n)}_{xy}=\mathbf{f}^{(n)}(x,y)-\bar{\mathbf{f}}(x,y)$ denote the deviation of the $n$-th Monte Carlo flow sample from the mean. The per-pixel covariance is then:
$\mathbf{\Sigma}(x,y)=\frac{1}{N_{s}}\sum_{n=1}^{N_{s}}\mathbf{d}^{(n)}_{xy}\left(\mathbf{d}^{(n)}_{xy}\right)^{\!\top}$ (9)
A value of $N_{s}=20$ Monte Carlo iterations was used, based on preliminary testing that showed no noticeable differences in the covariance matrices beyond this number of iterations. This yields a $2\times 2$ covariance matrix per pixel, encoding both the variance and correlation of the horizontal and vertical flow components. To maintain consistency with the raw event-based data, both the resulting flow fields and their associated covariance maps were distorted using the camera intrinsics before being stored in the dataset.
8 Sim-to-real Evaluation
Building on the real data acquisition, we now evaluate the capacity of models trained exclusively on the synthetic eStonefish-Scenes dataset to generalise to real underwater data. This evaluation aims to quantify the sim-to-real transfer performance and assess how well the proposed dataset supports learning event-based optical flow representations under real imaging conditions.
8.1 Network Architecture
The network architecture adopted in this work is inspired by the ConvGRU-based recurrent framework in (firenetHagenaars21). Similar to their design, our model is a time-driven convolutional recurrent network built around ConvGRU modules integrated within the encoder. A shallow convolutional head precedes the encoder, which consists of four downsampling blocks, each comprising a convolutional layer followed by group normalisation. The number of feature channels doubles after each down-sampling stage. The bottleneck section includes two residual blocks with skip connections. The decoder mirrors the encoder with four upsampling blocks that apply bilinear upsampling followed by a convolutional layer. Each upsampling stage includes a predictor block, composed of a single convolutional layer that outputs the optical flow at its respective resolution. A diagram of the network architecture is shown in Figure 9.
Figure 9: Architecture of the ConvGRU-based neural network used for optical flow prediction.
8.2 Dataset Preparation
For training, we employed the static sequences of the eStonefish-Scenes dataset. Dynamic sequences were excluded to avoid inconsistencies in the smoothness loss caused by independent object motion. For evaluation, we used the real-world dataset, described in Section 7, to assess sim-to-real generalisation. The input event data were encoded using the Gaussian temporal encoding scheme proposed by (encodingDing21), which preserves both spatial and temporal event distributions.
Input volumes were centre-cropped to $256\times 256$ pixels. Spatial augmentations comprised random horizontal and vertical flips. Temporal robustness was encouraged by injecting synthetic event noise: for each network input, 200 additional events were added to the event stream. Training samples were formed by segmenting each sequence between every pair of consecutive grayscale frames. The network therefore received the events occurring strictly within $t,t+\Delta t]$ for successive grayscale timestamps.
8.3 Multi-scale Loss & Metrics
The network was trained using the self-supervised loss formulation introduced by evflownetZhu18, combining a photometric reconstruction term with a spatial smoothness regulariser. The photometric component measures the intensity difference between the reference grayscale frame and the flow-warped target frame, while the smoothness term enforces continuity of the predicted flow field within local neighbourhoods. The latter assumption holds for static, rigid environments, which motivated the use of only static sequences for training. Training loss was computed using the Charbonnier penalty to mitigate the effect of outliers.
Network performance was evaluated using the Average Endpoint Error (AEE) and the Average Angular Error (AAE), following standard definitions in the optical flow literature. In benchmarks such as KITTI, a pixel is typically considered an outlier when its endpoint error exceeds three pixels (or 5% of the flow magnitude), and only the percentage of such outliers is reported. This fixed threshold assumes uniform reliability across all pixels, which is unsuitable for event-based data where uncertainty varies spatially. To account for this, each pixel in our evaluation was assigned a weight $w_{xy}$ derived from the inverse of the area of its uncertainty ellipse computed from the per-pixel covariance matrix. The weight was clipped such that $w_{xy}\leq 1$. The weighted AEE was then defined as:
$\mathrm{AEE}_{w}=\frac{\sum_{x,y}w_{xy}\sqrt{(u_{xy}-u^{gt}_{xy})^{2}+(v_{xy}-v^{gt}_{xy})^{2}}}{\sum_{x,y}w_{xy}}$ (10)
The weighted AAE was computed as:
$\mathrm{AAE}_{w}=\frac{\sum_{x,y}w_{xy}\arccos\left(\frac{u_{xy}u^{gt}_{xy}+v_{xy}v^{gt}_{xy}+1}{\sqrt{(u_{xy}^{2}+v_{xy}^{2}+1)({u^{gt}_{xy}}^{2}+{v^{gt}_{xy}}^{2}+1)}}\right)}{\sum_{x,y}w_{xy}}$ (11)
This formulation integrates the spatially varying confidence of the optical flow, attenuating the contribution of uncertain regions while preserving global consistency. It therefore provides a more representative measure of performance in real-world, noise-prone event-based scenarios.
8.4 Training and Evaluation
Our model was trained on an NVIDIA Quadro RTX 6000 GPU with 24 GB VRAM for 200 epochs with a batch size of 16. We used the AdamW optimiser with a weight decay of $0.01$ and learning rate of $0.004$, decayed using a polynomial learning rate scheduler with a warm-up of $3$ epochs. The models were implemented in PyTorch $2.9.1$ and Python $3.8$. The source code with all hyperparameter configurations and pre-trained models will be made available at  https://github.com/CIRS-Girona/ebof.
The model was then evaluated on a standard laptop equipped with an NVIDIA RTX 3070 GPU running Ubuntu 20.04, Python $3.8$ and PyTorch $2.8.1$. For validation, the network was tested on a held-out subset corresponding to 10% of the synthetic eStonefish-Scenes dataset, comprising 607 input samples, as well as on the real-world dataset, comprising 5,984 input samples. No fine-tuning was performed on the real data. Performance was assessed using the uncertainty-aware Average Endpoint Error (AEE) and Average Angular Error (AAE) metrics described in Section 8.3.
8.5 Results
8.5.1 Synthetic Data Evaluation
We first evaluate the performance of the proposed ConvGRU-based network on a held-out subset of the synthetic eStonefish-Scenes dataset. This subset corresponds to 10% of the available synthetic data and comprises 607 input samples, each formed by encoding the events occurring between two consecutive grayscale frames. The evaluation set includes both down-looking and front-looking camera configurations and covers a range of motion patterns, including forward and backward translation, vertical ascent and descent, and combined translational and rotational motion.
Figure 10 presents representative qualitative results from the synthetic test set. For each example, we show the reference grayscale frame, the encoded event input, the ground-truth optical flow, and the predicted flow. The first row illustrates a down-looking forward translation, where the predicted flow closely matches the ground truth in both magnitude and direction, achieving an endpoint error of $EE=0.329$ and an angular error of $AE=0.096$. The second and third rows demonstrate front-looking and down-looking descent motions, respectively. In these cases, the network accurately recovers the dominant vertical flow components induced by heave, although small discrepancies appear in regions with sparse events or low texture. The final example combines backward translation with rotational motion, producing more complex flow patterns. While the predicted flow captures the overall structure of the motion, local deviations are more pronounced. Across all examples, the model preserves coherent flow directions in regions with sufficient event activity. These qualitative results corroborate the quantitative findings and highlight the network’s ability to generalise across diverse synthetic underwater scenarios.
Figure 10: Qualitative results on the synthetic eStonefish-Scenes test set. Each row corresponds to a different camera configuration and motion pattern. Columns show (from left to right) the reference grayscale image, the encoded event input accumulated between two consecutive grayscale frames, the ground-truth optical flow, and the predicted optical flow produced by the ConvGRU-based network. Optical flow is visualised using a colour-coded representation with arrows indicating flow direction and magnitude. The Endpoint Error ($E$) and Angular Error (AE) for each example are reported on the right. The results demonstrate that the network accurately captures dominant translational motion and preserves coherent flow directions across diverse synthetic underwater scenarios.
8.5.2 Real Data Evaluation
Figure 11 presents qualitative results obtained on the real-world dataset. Each example shows the reference grayscale image, the encoded event input, the ground-truth optical flow computed via frame-to-poster registration, and the corresponding network prediction. Across all examples, the predicted flow fields exhibit strong qualitative agreement with the ground truth in regions with sufficient event activity, accurately capturing dominant motion directions and flow structure. Discrepancies are primarily observed near textureless regions, depth discontinuities, or areas affected by registration uncertainty, where the underlying ground truth is itself less reliable. Notably, the network preserves coherent flow patterns even under complex motion involving rotation, demonstrating robustness to real-world sensing artefacts despite being trained exclusively on synthetic data.
Figure 11: Qualitative results on the eStonefish-Scenes real-world validation dataset. Each row corresponds to a different ROV motion pattern acquired using a DAVIS346 event camera mounted on a BlueROV2 platform. Columns show (from left to right) the reference grayscale image, the encoded event input accumulated between consecutive grayscale frames, the ground-truth optical flow computed via frame-to-poster homography registration, and the predicted optical flow produced by the ConvGRU-based network. The results demonstrate strong qualitative agreement between prediction and ground truth across diverse real underwater scenarios, despite the network being trained exclusively on synthetic data.
8.5.3 Quantitative Evaluation
Dataset AEE (Pixels) AAE (Radians)
Synthetic eStonefish-Scenes (Test set, 10%) 0.636 0.173
Real eStonefish-Scenes 0.79 0.2
Table 2: Quantitative evaluation results on the synthetic and real-world eStonefish-Scenes datasets. AEE is measured in pixels and AAE in radians.
Table 2 reports quantitative performance on both the synthetic eStonefish-Scenes test set and the real-world dataset. For the synthetic data, where dense ground-truth optical flow is available from simulation, performance is evaluated using the unweighted AEE and AAE. Under these conditions, the proposed ConvGRU-based model achieves an average endpoint error of 0.636 pixels and an average angular error of 0.173 radians.
For the real-world dataset, where ground truth is obtained via homography-based registration and is inherently affected by calibration and feature-matching uncertainty, performance is evaluated using uncertainty-aware metrics. In this setting, the model attains an uncertainty-weighted endpoint error of 0.79 pixels and an uncertainty-weighted angular error of 0.2 radians without any fine-tuning on real data. The observed performance gap between synthetic and real data reflects the expected sim-to-real discrepancy induced by sensor noise, optical distortions, and imperfect ground-truth estimation. Importantly, incorporating uncertainty into the evaluation attenuates the influence of ambiguous regions and yields more reliable performance estimates on real underwater data.
@@ 9 Conclusion
This work addresses the challenge of developing reliable optical flow prediction networks for underwater environments, where collecting large-scale, annotated, event-based optical flow datasets is exceptionally difficult due to the complexities of underwater imaging. To overcome this, we introduce the eStonefish-Scenes dataset, along with a user-friendly pipeline to extend and adapt it. To the best of our knowledge, no prior event-based datasets exist for underwater applications. The dataset incorporates scenes capturing event-based data from an ROV navigating a dynamic, coral-rich environment filled with diverse marine flora and fauna. We also introduce Stonefish-SceneGen, an automated scene generator that eliminates the tedious and time-consuming process of manually creating reef environments. Additionally, we present Stonefish-Boids, a package that simulates realistic schooling behaviour of fish. The dataset and supporting packages were created with the goal of accelerating research on EBCs for underwater applications, especially for training CNNs and SNNs. We also present eWiz, a comprehensive event-based data manipulation library that includes an optimised event data format. The library features data converters, writers, readers, and processing modules focused on data management, compression, and optimisation. Additionally, it offers various loss functions, performance metrics, and visualisers to facilitate work with these datasets, along with an implementation of the contrast maximisation algorithm. Both eWiz and eStonefish-Scenes are designed to streamline research on EBCs and motion estimation for autonomous underwater vehicles by providing easy-to-use libraries for generating, processing, and evaluating event-based datasets. In the long run, eWiz will support not only optimisation-based algorithms but also deep learning-based and SNN-based alternatives. We plan to add more functionalities, such as ground truth generation via optimisation and SNN-based training algorithms.
Beyond dataset generation, we validated the practical utility of eStonefish-Scenes through a real-world experimental study using an event-based camera mounted on a BlueROV2 platform. Ground-truth optical flow was computed via homography-based frame-to-poster registration, and per-pixel uncertainty was explicitly estimated and incorporated into the evaluation metrics. A ConvGRU-based optical flow network trained exclusively on synthetic data demonstrated effective sim-to-real transfer when evaluated on real underwater sequences, with qualitative and quantitative results showing strong agreement in dominant motion patterns despite the absence of real-data fine-tuning. The combined quantitative and qualitative results demonstrate that the proposed synthetic dataset supports effective learning of transferable event-based motion representations for underwater environments. The use of uncertainty-aware evaluation metrics provides a more reliable assessment of performance in real-world settings, where ground truth is inherently imperfect. By downweighting regions with high uncertainty, the proposed evaluation framework avoids over-penalising predictions in ambiguous areas and offers a principled alternative to fixed outlier thresholds commonly used in optical flow benchmarks. These results suggest that synthetic event-based datasets, when combined with realistic simulation, can substantially reduce the reliance on large-scale real-world data collection for underwater robotics applications.
Future work will focus on extending the dataset to include dynamic scenes and independently moving objects, as well as exploring spiking neural network architectures to more fully exploit the temporal nature of event-based data.
@@
Jad Mansour was supported by the Joan Oró Grant no. 2024 FI-2 00762. This study was also supported in part by the Spanish government through the ASSIST project PID2023-149413OB-I00 and the IURBI project CNS2023-144688.
@@
Jad Mansour: Conceptualization, Methodology, Software, Data Curation, Writing - original draft; Sebastian Realpe: Software, Writing - original draft; Hayat Rajani: Conceptualization, Methodology, Supervision, Validation, Writing - original draft; Michele Grimaldi: Software, Writing - review & editing; Rafael Garcia: Methodology, Project Administration, Supervision, Writing - review & editing; Nuno Gracias: Project Administration, Supervision, Writing - review & editing.